OpenStudy (anonymous):

8 – 10x + x2, for x = 5

OpenStudy (ganpat):

substitute value of x as 5 in equation.. and evaluate..

OpenStudy (anonymous):

this is what igot 10*5=50 8-50=-42 5*2=10 -42+10=-32

OpenStudy (ganpat):

8 - 10 (5) + 5 sq 8 - 50 + 25 its actually, 17...

OpenStudy (ganpat):

its x sq... so 5 square... Not 5 * 2..

OpenStudy (anonymous):

how is it sq

OpenStudy (ganpat):

because, if it is would have been multiplication than, they wold write as 2x...